A box of mass 6.2 kg slides 4.8 m down a ramp that is inclined 38 degrees to the horizontal. What is the velocity of the box right before it hits the ground?
Answer #1

Here ,

let the speed of the box at the bottom of the incline is v

Using conservation of energy

0.50 * m * v^2 = m * g * L * sin(theta)

0.50 * v^2 = 9.8 * 4.8 * sin(38)

solving for v

v = 7.7 m/s

the velocity of the box right before hitting the ground is 7.7 m/s